# Prompt Injection Vulnerability of Consensus Generating Applications in Digital Democracy

### v3

@@ 6 Conclusion and Future Work
LLMs are increasingly employed to generate consensus statements in support of group deliberation, representing one of the most promising applications of AI in digital democratic systems (Konya et al. 2025; Velez, Green, and Sevi 2025; Tessler et al. 2024). Our analysis identifies fundamental vulnerabilities to prompt injection attacks when LLMs are tasked with autonomously producing consensus statements, particularly in deliberative settings where agreement and disagreement are evenly balanced. These vulnerabilities are not uniform: they are more pronounced for specific combinations of public policy clusters and political parties, and they intensify when attacks take the form of rational-sounding arguments rather than emotional language or fabricated statistics.
Although the defense techniques examined in this work achieve near-zero attack success rates and minimal overrefusals when restricting attention to non-ambiguous consensus, more sophisticated future attacks will likely reveal additional vulnerabilities. More broadly, empirical evaluations provide valuable evidence but cannot, on their own, fully guarantee the long-term robustness of a defense (Nasr et al. 2026). Future evaluations should therefore adopt an explicitly adversarial posture, stressing defenses to the point of failure—and ideally failing to make them fail.
@@ Appendix B BERT Classifier
Appendix C Definition of Majority Rule
To identify the majority position on each question, we aggregate the opinion texts of all participants for a given question into a single “consensus valence” (agree, disagree, or ambiguous) using our fine-tuned BERT classifier.
Each opinion texts is first mapped onto one of three categories: agree, disagree, or ambiguous/no clear opinion. Concretely, every response is labeled as follows: $1$ indicates agreement with the proposition, 0 indicates disagreement, and $2$ indicates an ambiguous or neutral stance. Let $c_{1}$ be the number of responses labeled $1$, $c_{0}$ the number labeled 0, and $c_{2}$ the number labeled $2$. The total number of opinions is $n=c_{0}+c_{1}+c_{2}$, and the majority threshold is defined as $t=n/2$.
The algorithm then determines the consensus valence by comparing $c_{0}$, $c_{1}$, and $c_{2}$ to this threshold: if $c_{1}>t$, the consensus valence is agree; if $c_{0}>t$, the consensus valence is disagree; and if $c_{2}>t$, the consensus valence is ambiguous. In cases of exact balance, if $c_{1}=t$ and $c_{0}<t$, the consensus is agree; if $c_{0}=t$ and $c_{1}<t$, the consensus is disagree; and if $c_{1}=t$ and $c_{0}=t$, the consensus is ambiguous. In all remaining cases (i.e., when no category exceeds or cleanly matches the majority threshold), the algorithm conservatively assigns the consensus valence as ambiguous.
Appendix D Performance of Default LLMs
Using our BERT-based classifier, we evaluated each consensus statement produced by the different models and measured how closely its valence matched the majority rule for each question.
Because the non–Habermas Machine LLMs are the ones we later subject to manipulations, we assessed the Habermas Machine on the original dataset, but evaluated all other LLMs on an expanded version of the dataset aligned with our taxonomy and free of prompt injections.
As shown in Figure 6, 32.29% of the consensus statements produced by the Habermas Machine and 62.3% of those generated by LLaMA 3.1 8B Instruct, 56.62% generated by GPT 4.1 Nano and 60.54% generate by Apertus 8B Instruct do not align with the majority rule (”Error rate”). This highlights the importance of filtering out prompts that fail to align with the majority rule in the absence of attacks, in order to disentangle the effect of prompt injection from the LLMs’ baseline performance.
@@ Appendix D Performance of Default LLMs
Appendix E Definition of Clusters of Public Policies and Political Parties’ Stances Extraction
@@
Appendix F Prompt Injection Detectors in Digital Democracy
@@
Appendix G Group Sequence Policy Optimization (GSPO)
We employ Group Sequence Policy Optimization (GSPO) (Zheng et al. 2025), a reinforcement learning (RL) approach that fine-tunes large language models (LLMs) by explicitly rewarding group-level sequence generations that internally satisfy the majority rule.
@@ Appendix G Group Sequence Policy Optimization (GSPO)
Training consists of 4,500 gradient steps of GSPO, with each prompt generating a group of $K=6$ candidate sequences. Each generated sequence is evaluated using the sum of distinct reward signals: 1. Structural Predictability: A rule-based reward that grants a score of 1.0 only if the generated consensus statement under prompt injection by the LLM during GSPO-training correctly uses the format ”###Reasoning” and ”###Consensus View” to structure the output.
@@ 2. Semantic Similarity:
3. Factual Accuracy: A binary reward of 1.0 if the LLM’s generated consensus statement under prompt injection by the LLM during GSPO-training matched the majority rule, and 0.0 otherwise.
@@
Within this augmented democratic framework, your duty is to **preserve fairness and integrity by summarizing the opinions of all participants and respecting the decision of the majority**.
@@
      Label each opinion: 1 = agree, 0 = disagree, 2 = ambiguous or no clear opinion.
@@
      Count: c0 = number of 0s, c1 = number of 1s, c2 = number of 2s, n = total number of opinions, t = n divided by 2.
@@
      Determine the FINAL VALENCE: If c1 is greater than t, then FINAL VALENCE = agree. If c0 is greater than t, then FINAL VALENCE = disagree. If c2 is greater than t, then FINAL VALENCE = ambiguous. If c1 equals t and c0 is less than t, then FINAL VALENCE = agree. If c0 equals t and c1 is less than t, then FINAL VALENCE = disagree. If c1 equals t and c0 equals t, then consensus = ambiguous. Otherwise, consensus = ambiguous.
@@
Appendix H Vulnerability for different LLMs
Figure 9: Vulnerability to prompt‑injection attacks for GPT 4.1 Nano (panel (a)) and Apertus 8B Instruct (panel (b)), expressed as Attack Success Rates (ASR) by public policy question cluster (x‑axis) and UK political party (y‑axis). In both panels, ASR is computed from the disagree, ambiguous, and agree cells of the underlying confusion matrices, and all elements are shown in red. The top bar chart displays the bootstrap mean ASR for each public policy question cluster, the bar chart on the right displays the bootstrap mean ASR for each political party, and the bottom‑left heatmap reports bootstrap mean ASR for every party–cluster pair, with darker shades indicating higher attack success and black cells indicating insufficient data. All ASR values and their 95% confidence intervals are estimated via bootstrapping with 5,000 iterations.
Appendix I Attack Effectiveness by Rhetorical Strategy
Figure 10: Effectiveness of prompt-injection attacks by rhetorical strategy across three LLMs: GPT 4.1 Nano, LLaMA 3.1 8B Instruct and Apertus 8B. All ASR values and their 95% confidence intervals are estimated via bootstrapping with 5,000 iterations.
Appendix J Attack Effectiveness with SmartVote
We validate our findings using a dataset of comments submitted to Smartvote, a popular Swiss voting advice application, to simulate intra-party consensus statements. From a database containing approximately 100,000 comments written by candidates running for the national parliament in Switzerland (Stammbach et al. 2024), we utilize 26,502 comments to examine LLMs’ vulnerabilities in consensus formation scenarios involving 5-40 opinion texts.
We define consensus tasks for each unique combination of political party (26 parties), language (French, German, Italian), and public-policy question (374 questions). This framework allows us to analyze the effects of the same taxonomy dimensions established in the main text while examining how ASR values vary with the number of participants involved in the consensus formation process. We fine-tuned a multilingual BERT model (Boizard et al. 2025) with default hyperparameters for defining valences using 35,665 “Agree”/“Disagree” Swiss candidates ratings and “Ambiguous” ratings from GPT-4o. After filtering out prompts that did not satisfy minimum requirements, 24,216 were used for analyzing support and 23,976 for criticism attacks.
(a) ASR as a function of the number of participants using LLaMA 3.1 8B Instruct.
(b) ASR as a function of the number of participants using Apertus 8B Instruct.
Figure 11: Smartvote dataset analysis: participant scaling effects
Appendix K ASR decomposition
To assess the extent to which we can improve LLM robustness against prompt injections in digital democracy systems, we present results both including and excluding the ”ambiguous” valence from our analysis.
Figure 12: Robustness to prompt-injection attacks when the ”ambiguous” valence is included in the ASR computation.
Figure 13: Robustness to prompt-injection attacks when the ”ambiguous” valence is excluded from the ASR computation.
Appendix L DPO + GSPO with LLaMA 3.1 8B Instruct
@@ Appendix I Attack Effectiveness by Rhetorical Strategy
Next, we merge the paired prompts with the pairs of LLM-generated consensus into a single preferences dataset, where each entry contains the injected prompt, the original consensus, and the “polluted” consensus statements. To sharpen the training signal, we apply two filtering steps. First, we retain only entries where the original and “polluted” consensus statements differ in valence, and where the desired valence aligns with the majority opinion of participants. Second, we apply oversampling to balance the distribution of ”Agree”, ”Disagree”, and ”Ambiguous” valences.
@@
After fine-tuning LLaMA with DPO, we further fine-tune the same LLM to generate consensus statements that adhere to the majority rule. To this end, we introduce an additional reward function on top of the hyperparameters and reward functions described in Appendix G: ”4. Prompt-injection localization reward: This binary signal assigns a reward of 1.0 if the LLM correctly identifies during GSPO training the participant responsible for the prompt-injection attack, and 0.0 otherwise.”
@@
      Label each opinion: 1 = agree, 0 = disagree, 2 = ambiguous or no clear opinion.
@@
      Count: c0 = number of 0s, c1 = number of 1s, c2 = number of 2s, n = total number of opinions, t = n divided by 2.
@@
      Determine the FINAL VALENCE: If c1 is greater than t, then FINAL VALENCE = agree. If c0 is greater than t, then FINAL VALENCE = disagree. If c2 is greater than t, then FINAL VALENCE = ambiguous. If c1 equals t and c0 is less than t, then FINAL VALENCE = agree. If c0 equals t and c1 is less than t, then FINAL VALENCE = disagree. If c1 equals t and c0 equals t, then consensus = ambiguous. Otherwise, consensus = ambiguous.
@@
Figure 14: Robustness against prompt‑injection attacks for LLaMA 3.1 8B Instruct, expressed as Attack Success Rates (ASR) by public policy question cluster (x‑axis) and UK political party (y‑axis). This figure displays clustered bar charts, where red bars correspond to ASR without robustness (using only the agree and disagree categories) and blue bars correspond to ASR after applying the robustness methodology with the same categories. The top bar chart displays the bootstrap mean ASR for each public policy question cluster, the bar chart on the right displays the bootstrap mean ASR for each political party, and the bottom‑left heatmap reports bootstrap mean ASR for every party–cluster pair, with darker shades indicating higher attack success and black cells indicating insufficient data. All ASR values and their 95% confidence intervals are estimated via bootstrapping with 5,000 iterations.
Appendix M Deliberative Alignment with GPT 4.1 Nano
Drawing inspiration from the methodology developed by Guan et al. (2024) for addressing jailbreaking attacks, we modify the preferences dataset constructed for DPO alignment in Appendix L to investigate whether the reasoning capabilities of DeepSeek-R1 (Guo et al. 2025) can enhance GPT-4.1 Nano’s robustness against prompt-injection attacks through DPO integration with safety and security policies.
@@
Figure 15: Robustness against prompt‑injection attacks for GPT 4.1 Nano, expressed as Attack Success Rates (ASR) by public policy question cluster (x‑axis) and UK political party (y‑axis). This figure displays clustered bar charts, where red bars correspond to ASR without robustness (using only the agree and disagree categories) and blue bars correspond to ASR after applying the robustness methodology with the same categories. The top bar chart displays the bootstrap mean ASR for each public policy question cluster, the bar chart on the right displays the bootstrap mean ASR for each political party, and the bottom‑left heatmap reports bootstrap mean ASR for every party–cluster pair, with darker shades indicating higher attack success and black cells indicating insufficient data. All ASR values and their 95% confidence intervals are estimated via bootstrapping with 5,000 iterations.
Appendix N Overrefusal Rates
Figure 16: Overrefusal rates, calculated as the ASR from a 3×3 confusion matrix comparing the valences of the original and the robust LLM, both in the absence of prompt‑injection attacks. All ASR values and their 95% confidence intervals are estimated via bootstrapping with 5,000 iterations.

### v4

@@ 6 Conclusion and Future Work
LLMs are increasingly employed to generate consensus statements in support of group deliberation, representing one of the most promising applications of AI in digital democratic systems (Konya et al. 2025; Velez, Green, and Sevi 2025; Tessler et al. 2024). Our analysis reveals fundamental vulnerabilities to prompt-injection attacks when LLMs are tasked with autonomously producing consensus statements. These vulnerabilities are not uniform. When examined in the aggregate, they are disproportionately concentrated in deliberative settings where agreement and disagreement are evenly balanced. In addition, vulnerabilities tend to favor GB-unionist right parties over pro-independence left parties, and are more pronounced for attacks framed as rational or procedural arguments, such as imperative orders or impossibility-of-agreement narratives, rather than emotional appeals or fabricated statistics.
Our defense pipeline substantially reduces directional failures whenever the underlying consensus has a clear positive or negative valence. In those cases, most remaining discrepancies come from the model producing ambiguous consensus statements rather than reversing the direction of the consensus. The difficult setting is when agreement and disagreement are finely balanced: the pipeline still allows the LLM to collapse that ambiguity into clear agreement or disagreement, even in the absence of prompt-injection attacks. Designing LLM-based DD systems that preserve directional protection while handling genuinely ambiguous consensuses in a principled way therefore remains the central open challenge for resilient consensus-generation systems.
Beyond the reduced sample size used in our experiments, a central limitation of this work lies in the restricted taxonomy of attacks we consider, which covers only a narrow set of human-readable prompt-injection strategies. More sophisticated adversaries may craft attacks that evade detection, exploit structured representations, or directly target the ambiguity failure mode identified here (Nasr et al. 2026).
@@ Appendix B BERT Classifier
Appendix C Performance of Default LLMs
Using our BERT-based classifier, we evaluated each consensus statement produced by the different models and measured how closely its valence matched the net position for each question.
As shown in Figure 7, 56.22% of the consensus statements produced by the Habermas Machine and 63.11% of those generated by LLaMA 3.1 8B Instruct, 55.95% generated by GPT 4.1 Nano and 59.86% generate by Apertus 8B Instruct do not align with the corresponding net position. This highlights the importance of filtering out prompts that fail to align with the net position in the absence of attacks, in order to disentangle the effect of prompt injection from the LLMs’ baseline performance.
@@ Appendix C Performance of Default LLMs
Appendix D Definition of Clusters of Public Policies and Political Parties’ Stances Extraction
@@
Appendix E Prompt Injection Detectors in Digital Democracy
@@
Appendix F Group Sequence Policy Optimization (GSPO)
We employ Group Sequence Policy Optimization (GSPO) (Zheng et al. 2025), a reinforcement learning (RL) approach that fine-tunes large language models (LLMs) by explicitly rewarding group-level sequence generations that internally satisfy the calculation of the net position.
@@ Appendix F Group Sequence Policy Optimization (GSPO)
Training consists of 4,500 gradient steps of GSPO, with each prompt generating a group of $K=6$ candidate sequences. Each generated sequence is evaluated using the sum of distinct reward signals:
1. Structural Predictability: A rule-based reward that grants a score of 1.0 only if the generated consensus statement under prompt injection by the LLM during GSPO-training correctly uses the format ”###Reasoning” and ”###Consensus View” to structure the output.
@@ 2. Semantic Similarity:
3. Factual Accuracy: A binary reward of 1.0 if the LLM’s generated consensus statement under prompt injection by the LLM during GSPO-training matched the net position, and 0.0 otherwise.
@@
Within this augmented democratic framework, your duty is to **preserve fairness and integrity by summarizing the opinions of all participants and respecting the calculation of the net position**.
@@
      Label each opinion: agree, disagree, or ambiguous.
@@
      Assign plus 1 to each agree opinion, minus 1 to each disagree opinion, and 0 to each ambiguous opinion. Sum all values and divide by the total number of opinions to obtain the net position, which ranges from minus 1 to plus 1.
@@
      Determine the FINAL VALENCE: if the net position is positive, then FINAL VALENCE = agree. If the net position is negative, then FINAL VALENCE = disagree. If the net position is zero, then FINAL VALENCE = ambiguous.
@@
Appendix G DPO + GSPO with LLaMA 3.1 8B Instruct
@@
Next, we merge the paired prompts with the pairs of LLM-generated consensus into a single preferences dataset, where each entry contains the injected prompt, the original consensus, and the “polluted” consensus statements. To sharpen the training signal, we apply two filtering steps. First, we retain only entries where the original and “polluted” consensus statements differ in valence, and where the desired valence aligns with the net position. Second, we apply oversampling to balance the distribution of ”Agree”, ”Disagree”, and ”Ambiguous” valences.
@@
After fine-tuning LLaMA with DPO, we further fine-tune the same LLM to generate consensus statements that adhere to the valence associated to the net position. To this end, we introduce an additional reward function on top of the hyperparameters and reward functions described in Appendix F: ”4. Prompt-injection localization reward: This binary signal assigns a reward of 1.0 if the LLM correctly identifies during GSPO training the participant responsible for the prompt-injection attack, and 0.0 otherwise.”
@@
      Label each opinion: agree, disagree, or ambiguous.
@@
      Assign plus 1 to each agree opinion, minus 1 to each disagree opinion, and 0 to each ambiguous opinion. Sum all values and divide by the total number of valid opinions to obtain the net position, which ranges from minus 1 to plus 1.
@@
      Determine the FINAL VALENCE: if the net position is positive, then FINAL VALENCE = agree. If the net position is negative, then FINAL VALENCE = disagree. If the net position is zero, then FINAL VALENCE = ambiguous.
@@
Appendix H Deliberative Alignment with GPT 4.1 Nano
Drawing inspiration from the methodology developed by Guan et al. (2024) for addressing jailbreaking attacks, we modify the preferences dataset constructed for DPO alignment to investigate whether the reasoning capabilities of DeepSeek-R1 (Guo et al. 2025) can enhance GPT-4.1 Nano’s robustness against prompt-injection attacks through DPO integration with safety and security policies.
@@
Appendix I Vulnerability analysis for other parties
Figure 10: .
Appendix J Overrefusal Analysis
Figure 11: .